OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y x - 5 y -x - 4

OpenStudy (e.mccormick):

Looks like the question is incomplete. However, the basic method is to solve the system and then you know where the lines intersect. If this is actually: \(y=x-5\) \(y=-x-4\) You can do this: \(x-4=x-5\) Then solve for x, put x into either of the first equations and solve for y.